OpenStudy (onepieceftw):

Can someone check my work? How many grams of iron metal do you expect to be produced when 305 grams of a 75.5 percent by mass iron(II) nitrate solution reacts with excess aluminum metal? Show all of the work needed to solve this problem. 2 Al (s) + 3 Fe(NO3)2 (aq) yields 3 Fe (s) + 2 Al(NO3)2 (aq)

OpenStudy (onepieceftw):

305 g Fe(NO3)2 * 75.5 g Fe(NO3)2/100 g Fe(NO3)2 * 1 mol Fe(NO3)2 = 1.26 mol Fe(NO3)2 1.26 mol Fe(NO3)2 * 3 mol Fe/2 mol Al = 1.89 1.89 * 55.845 g Fe/1 mol Fe = 105.55 g Fe

OpenStudy (aaronq):

The molar ratio you used in step 2 is wrong. no need to use aluminum

OpenStudy (onepieceftw):

Would I use Iron (II) nitrate? Which would make the answer 70.36

OpenStudy (aaronq):

i'm not sure what you're referring to

OpenStudy (aaronq):

your error is in step 2 1.26 mol Fe(NO3)2 * 3 mol Fe/2 mol Al = 1.89

OpenStudy (onepieceftw):

The mole ratio, what would I use to determine the moles of Iron metal?

OpenStudy (aaronq):

yep that would be correct.

OpenStudy (onepieceftw):

Thanks bud.

OpenStudy (aaronq):

np